OpenStudy (anonymous):

Linear Algebra and Modular Math: We know that if a matrix A is n x n, then Ax = b (x and b are vectors) has a unique solution iff A is invertible. 2 questions: 1) If A is singular, then are there many solutions to Ax = b? Or are there no solutions? 2) If we are working with modular math and "=" is the congruence relation, then does the first statement I wrote still hold? Or should there be any changes in that statement?

ganeshie8 (ganeshie8):

familiar with subspaces ?

OpenStudy (anonymous):

Vaguely.. I have taken linear algebra about 3 years ago and returning to it now since I'm taking cryptography haha.

ganeshie8 (ganeshie8):

the linear combinations of column vectors form column space when A is singular, there will be many points(b's) that lie outside the column space, so you won't be able to reach all points by taking the combinations of column vectors

ganeshie8 (ganeshie8):

Consider below system : x + y = 2 2x+2y = 4 A : 1 1 2 2 b: 2 4

ganeshie8 (ganeshie8):

above system has a solution because the vector \(\large \binom{2}{4}\) is in column space of A

ganeshie8 (ganeshie8):

For the same matrix, A, there will NOT be a solution for a random vector like, \(\large \binom{1}{3}\) : \[\large \begin{pmatrix} 1&1\\2&2 \end{pmatrix}\begin{pmatrix} x\\y\end{pmatrix} = \begin{pmatrix} 1\\3\end{pmatrix}\]

ganeshie8 (ganeshie8):

To answer your specific question : 1) If A is singular, then are there many solutions to Ax = b? Or are there no solutions? Ax=b has infinitely many solutions when `b` is in column space of A Ax=b has 0 solutions otherwise

OpenStudy (anonymous):

Ah, thank you for that, I could see why from your explanations. Thank you!

ganeshie8 (ganeshie8):

Could you shed some more light on what you mean by \(\large Ax \equiv b\) ?

OpenStudy (anonymous):

So the second question.. what would change from all of these if Ax=b is under mod n?

ganeshie8 (ganeshie8):

\(\large Ax\equiv b\pmod n\) ? what exactly is this ?

OpenStudy (anonymous):

Oh, I mean for example, in regular (sorry for the lack of a better term) math, a matrix A is invertible if the determinant is not 0, but in modular math, an added condition is that A and n (if we're working in mod n) must be relatively prime. I was wondering if there's also a similar added condition in this context.

ganeshie8 (ganeshie8):

Hmm... let me try and see if I get what you're doing. when \(a\) is a scalar, \(\large ax\equiv 1\mod n\) has a solution, right ?

ganeshie8 (ganeshie8):

how do you define congruence in context of matrix ? do you take mod for each element ?

OpenStudy (anonymous):

Hmm maybe I should not have said congruence.. maybe that was careless writing on my part. I was really just wondering what would happen when the numbers in the matrix and vectors in Ax=b are allowed to have equivalent numbers mod n.

OpenStudy (anonymous):

anyway, thank you very much! I'm about to eat dinner for now, so I'll be off the computer for a while

ganeshie8 (ganeshie8):

\[\large \begin{pmatrix} a_1&b_1\\a_2&b_2\end{pmatrix} \begin{pmatrix} x\\y\end{pmatrix}= \begin{pmatrix} c_1\\c_2\end{pmatrix} \] Versus \[\large \begin{pmatrix} a_1\mod n&b_1 \mod n\\a_2\mod n&b_2\mod n\end{pmatrix} \begin{pmatrix} x\\y\end{pmatrix}= \begin{pmatrix} c_1\mod n\\c_2\mod n\end{pmatrix} \]

ganeshie8 (ganeshie8):

what good is that hmm

OpenStudy (anonymous):

Yes. Would anything that we've said so far change if that's the case?

OpenStudy (anonymous):

It's okay haha. Anyway I'm really going now. These are just curious questions that I've thought of to solve the problems in my homework

ganeshie8 (ganeshie8):

\(\large (a_ix+b_iy) \mod n = c_i \mod n\)

ganeshie8 (ganeshie8):

I don't see any problem with the system as you're taking mod both sides

OpenStudy (anonymous):

Essentially the whole point of me to ask these questions is to figure out if two plaintexts can map to the same ciphertext in a Hill Cipher

ganeshie8 (ganeshie8):

Ahh I see :) i have no clue but it looks interesting It seems, to encrypt vector x, you're taking \(\large Ax \pmod{26}\)

ganeshie8 (ganeshie8):

Since \(A\) is invertible, there will be one-to-one relationship between `x` and `Ax` we need to think a bit more on what happens after taking mod 26

OpenStudy (anonymous):

@ganeshie8 Yes! That's exactly what I'm getting at. Sorry for beating around the bush so much.

ganeshie8 (ganeshie8):

\[\large Ax \equiv b \pmod n \implies x = A^{-1}b \pmod{n}\]

ganeshie8 (ganeshie8):

for \(\large A^{-1}\) to exist, \(\large \det(A)\) must not be \(0\) also the multiplicative inverse of \(\large \det(A)\) must exist in \(\large \mod n\)

ganeshie8 (ganeshie8):

I still don't see why/how you think there is a possibility for two plaintexts to map to same cipher text

OpenStudy (anonymous):

If the user of the cipher decides to use a matrix key that is singular.. then for certain ciphertexts, there could be two plaintexts that map to it, no? just like how Ax=b has infinitely many solutions when b is in the column space of A?